# Regular Black Holes from Proper-Time flow in Quantum Gravity and their Quasinormal modes, Shadow and Hawking radiation

### v1

@@ 1 Introduction
The typical approach to extend the domain of a classical field theory into the quantum realm is to reformulate it as a quantum field theory (QFT). Following this idea, Weinberg proposed that the effective QFT of gravity might possess an UV attractive non-Gaussian fixed point (NGFP), a condition called asymptotic safety (AS) [3]. Within the framework of non-perturbative renormalization group (RG) [4, 5, 6, 7], this existence was confirmed for the 4-dimensional case, despite the absence of a complete description, motivating the study of its implications for BH physics. In this work, we specifically adopt the proper time (PT) formulation of the flow equation, which has a possible interpretation as a coarse graining of a Wilsonian effective action, with the direct physical meaning of its characteristic length scale as the lattice spacing of the effective field theory (EFT) [8, 9]. Proper time functional flow equations have attracted considerable interest in recent years due to their effectiveness in addressing non-perturbative phenomena [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Moreover, they offer the advantage of preserving certain symmetries of the underlying action, making them a valuable tool in theoretical analyses [23, 24, 25].
@@
Quasinormal modes (QNMs) of black holes are characteristic oscillations that dominate the response of a black hole to perturbations [29, 30, 31, 32]. They carry unique signatures of the black hole’s mass, charge, angular momentum and parameters of the theory, making them essential tools for probing black hole properties and the underlying gravitational theory. In gravitational wave astronomy, QNMs are crucial for interpreting the ringdown phase of black hole mergers observed by detectors like LIGO and Virgo [33, 34]. Since QNMs are purely determined by the background geometry, they provide a direct test of general relativity in the strong-field regime. Additionally, studying QNMs in alternative gravity theories or in the presence of quantum corrections can reveal potential deviations from classical black hole predictions. Therefore, a great number of recent publication were devoted to calculations of quasinormal modes of various quantum corrected black holes [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52].
@@
The paper is organized as follows: In Sec. 1, we introduce the regular black hole geometry arising from an asymptotically safe gravity scenario. In Sec. 2, we discuss the perturbation equations and the structure of the effective potential. Sec. 3 presents the numerical methods and the resulting quasinormal spectra. Section 5 is devoted to analysis of particle motion and shadows cast by such quantum corrected black holes. In Sec. 4 we analyze the grey-body factors and Hawking radiation for photons. We conclude with a discussion of the physical implications of our findings and possible directions for future work.
@@ 1.2 Running gravitational constant from Asymptotic Safety
Here, the smearing function $\rho_{\Lambda}[\phi_{0}]$ encodes the blocking procedure that characterizes the RG transformation.
@@
Here, $\tilde{S}_{\Lambda}^{(2)}$ is the hessian of the gauge fixed Wilsonian effective action and of the ghost action, and $r_{\Lambda}(s)$ contains the regulating function. The flow equation can be treated with different approaches.
To regularize the theory, we utilize a family of schemes, parametrized by $m$ and $\varepsilon$. With a diagonal hessian, where $A_{\Lambda,i}$ are the coefficients of the laplacians, it can be expressed in the parametric form
$r_{\Lambda}(A_{\Lambda,i}s;m)=\frac{2}{\Gamma(m)}\left(1+\varepsilon\frac{\Lambda}{2}\partial_{\Lambda}\log A_{\Lambda,i}\right)\left(mA_{\Lambda,i}s\Lambda^{2}\right)^{m}e^{-mA_{\Lambda,i}s\Lambda^{2}}.$ (1.12)
In the following, we refer to two possible values of $\varepsilon$, using the same nomenclature as in [22]. For $\varepsilon=0$, we obtain the “C scheme”, while for $\varepsilon=1$ the “B scheme”.
@@ 3.4 Numerical results for quasinormal modes
In order to obtain accurate values of the quasinormal frequencies, including higher overtones, we employed the Leaver method, which requires that the coefficients in the corresponding differential equation be given in a rational form. To meet this requirement, we expressed the metric functions as a series expansion in powers of the small parameter $q$. The expansion was carried out up to order $q^{-10}$, which proved sufficient for computing the first five overtones, as demonstrated in Table 7. As seen in the table, for a near-extremal black hole, the frequency of the fourth overtone varies by less than $0.1\%$ when higher-order corrections are included, while the fundamental mode remains unchanged to within six decimal places.
@@ 4 Grey-body factors and Hawking radiation
It is also worth noting that primordial black holes (PBHs) have been considered as viable dark matter candidates, since they are non-luminous, non-relativistic, and could have formed in the early Universe independently of baryonic physics. In the standard Schwarzschild case, PBHs with initial masses $M>10^{17}g$ are typically considered viable dark matter candidates, as they would not have emitted enough Hawking radiation to be excluded by existing observational constraints. This translates into an upper bound on the fraction $f_{\text{PBH}}$ of dark matter that PBHs can constitute, depending on their mass. In our analysis, we have shown that regular black holes radiate less efficiently (figure 11) due to their lower temperature (figure 10), which qualitatively suggests that the lower bound on their allowed mass could be relaxed. As a consequence, the constraints on $M$ and $f_{\text{PBH}}$ are less stringent, potentially allowing a larger fraction of dark matter to be composed of PBHs. This behavior is consistent with the recently findings [120, 121, 122], who demonstrated that other regular PBHs can evade gamma-ray constraints more easily and thus expand the viable dark matter mass window.
@@ Acknowledgments
We would like to acknowledge A.Zhidenko for useful discussions. The work of A. S. was supported by University of Catania and INFN Sezione di Catania. A.S. is grateful for the hospitality of the Institute of Physics at Silesian University.

### v2

@@ 1 Introduction
The typical approach to extend the domain of a classical field theory into the quantum realm is to reformulate it as a quantum field theory (QFT). Following this idea, Weinberg proposed that the effective QFT of gravity might possess an UV attractive non-Gaussian fixed point (NGFP), a condition called asymptotic safety (AS) [3]. Within the framework of non-perturbative renormalization group (RG) [4, 5, 6, 7], this existence was confirmed for the 4-dimensional case, despite the absence of a complete description, motivating the study of its implications for BH physics. In this work, we specifically adopt the proper time (PT) formulation of the flow equation, which has a possible interpretation as a coarse graining of a Wilsonian effective action, with the direct physical meaning of its characteristic length scale as the lattice spacing of the effective field theory (EFT) [8, 9]. Proper time functional flow equations have attracted considerable interest in recent years due to their effectiveness in addressing non-perturbative phenomena [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Moreover, they offer the advantage of preserving certain symmetries of the underlying action, making them a valuable tool in theoretical analyses [23, 24, 25, 26].
@@
Quasinormal modes (QNMs) of black holes are characteristic oscillations that dominate the response of a black hole to perturbations [30, 31, 32, 33]. They carry unique signatures of the black hole’s mass, charge, angular momentum and parameters of the theory, making them essential tools for probing black hole properties and the underlying gravitational theory. In gravitational wave astronomy, QNMs are crucial for interpreting the ringdown phase of black hole mergers observed by detectors like LIGO and Virgo [34, 35]. Since QNMs are purely determined by the background geometry, they provide a direct test of general relativity in the strong-field regime. Additionally, studying QNMs in alternative gravity theories or in the presence of quantum corrections can reveal potential deviations from classical black hole predictions. Therefore, a great number of recent publication were devoted to calculations of quasinormal modes of various quantum corrected black holes [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55].
@@
The paper is organized as follows: In Sec. LABEL:sec:geometry, we introduce the regular black hole geometry arising from an asymptotically safe gravity scenario. In Sec. 2, we discuss the perturbation equations and the structure of the effective potential. Sec. 3 presents the numerical methods and the resulting quasinormal spectra. Section 5 is devoted to analysis of particle motion and shadows cast by such quantum corrected black holes. In Sec. 4 we analyze the grey-body factors and Hawking radiation for photons. We conclude with a discussion of the physical implications of our findings and possible directions for future work. Throughout the paper we adopt natural units $c=\hbar=1$.
@@ 1.2 Running gravitational constant from Asymptotic Safety
The blocking procedure, encoded in the smearing function $\rho_{\Lambda}[\phi_{0}]$, is at the core of the RG transformation. Its role is to zoom out the system, much like our eyes do when looking out of the window of a departing airplane: as we move farther from the ground, the details of the landscape gradually fade and merge with nearby structures. Similarly, in a spin lattice where each cell has a spin-up value of 1 and a spin-down value of -1, it consists of grouping spins into blocks and coarse-graining them into new effective spins. In this way, the details of the microscopic system fade out, but the essential information is maintained. This coarse-graining is not unique; in fact, we can take the effective spin as the average value of the spins inside the block, or alternatively, we can take the most frequent spin value. Non-universal quantities, such as the position of possible fixed points, are sensitive to different methods. On the other hand, universal quantities, including the critical exponents, remain unaffected.
In momentum space, the blocking procedure retains only the modes below the cutoff scale, while averaging out the UV modes through the regularization.
@@
Here, $\tilde{S}_{\Lambda}^{(2)}$ is the hessian of the gauge fixed Wilsonian effective action and of the ghost action, and the regulating function is contained in
$r_{\Lambda}(s)=\Lambda\partial_{\Lambda}\rho_{k,\Lambda}(s).$ (1.12)
The flow equation can be treated with different approaches. The coarse-graining procedure is not unique even in momentum space, due to the different regularizing functions we can choose, which encode different ways of integrating the UV modes. To analyze how the regularization may affect the observables, it is important to test the theory using different choices. For this purpose, we utilize a family of schemes, parametrized by $m$
$\rho_{k,\Lambda}(s;m)=\frac{\Gamma\left(m,msk^{2}\right)-\Gamma\left(m,ms\Lambda^{2}\right)}{\Gamma\left(m\right)}$ (1.13)
that implies
$r_{\Lambda}(s;m)=\frac{2}{\Gamma\left(m\right)}\left(ms\Lambda^{2}\right)^{m}e^{-ms\Lambda^{2}}.$ (1.14)
To guarantee the correct suppression of the modes, this regulator must be adapted to hessians where the laplacians have coefficients $A_{\Lambda,i}$. In this case we simply rescale the regulating functions by $A_{\Lambda,i}$. This rescaling can be applied in two different ways. First, if we rescale directly $r_{\Lambda}$, it results
$r_{\Lambda}(A_{\Lambda,i}s;m)=\frac{2}{\Gamma(m)}\left(mA_{\Lambda,i}s\Lambda^{2}\right)^{m}e^{-mA_{\Lambda,i}s\Lambda^{2}}.$ (1.15)
Alternatively, if we rescale $\rho_{k,\Lambda}$ as in
$\rho_{k,\Lambda}(A_{\Lambda,i}s;m)=\frac{\Gamma\left(m,mA_{\Lambda,i}sk^{2}\right)-\Gamma\left(m,mA_{\Lambda,i}s\Lambda^{2}\right)}{\Gamma\left(m\right)},$ (1.16)
the eq. (1.12) implies
$r_{\Lambda}(A_{\Lambda,i}s;m)=\frac{2}{\Gamma(m)}\left(1+\frac{\Lambda}{2}\partial_{\Lambda}\log A_{\Lambda,i}\right)\left(mA_{\Lambda,i}s\Lambda^{2}\right)^{m}e^{-mA_{\Lambda,i}s\Lambda^{2}}.$ (1.17)
We combine these two families into one, introducing a second parameter $\varepsilon$
$r_{\Lambda}(A_{\Lambda,i}s;m,\varepsilon)=\frac{2}{\Gamma(m)}\left(1+\varepsilon\frac{\Lambda}{2}\partial_{\Lambda}\log A_{\Lambda,i}\right)\left(mA_{\Lambda,i}s\Lambda^{2}\right)^{m}e^{-mA_{\Lambda,i}s\Lambda^{2}}.$ (1.18)
In the following, we use the same nomenclature as in [22]. For $\varepsilon=0$, the derivative of $\log A_{\Lambda,i}$ is absent, leading to the “C scheme”, while for $\varepsilon=1$ the derivative is present, corresponding to the “B scheme”.
@@ 3.2 Leaver method
The original form of the metric function is not rational, which prevents a direct application of the Leaver method. To overcome this, we expand the metric function in powers of the small parameter $q$ and retain enough terms so that the results of the calculations remain stable within the desired accuracy. In practice, a 10th-order expansion is usually sufficient, and we have verified this by extending the series up to the 16th order, finding no significant changes. A similar approach has been recently applied to the Bardeen black holes in [96].
@@ 3.4 Numerical results for quasinormal modes
In order to obtain accurate values of the quasinormal frequencies, including higher overtones, we employed the Leaver method, which requires that the coefficients in the corresponding differential equation be given in a rational form. To meet this requirement, we expressed the metric functions as a series expansion in powers of the small parameter $1/q$. The expansion was carried out up to order $q^{-10}$, which proved sufficient for computing the first five overtones, as demonstrated in Table 7. As seen in the table, for a near-extremal black hole, the frequency of the fourth overtone varies by less than $0.1\%$ when higher-order corrections are included, while the fundamental mode remains unchanged to within six decimal places.
@@ 4 Grey-body factors and Hawking radiation
It is also worth noting that primordial black holes (PBHs) have been considered as viable dark matter candidates, since they are non-luminous, non-relativistic, and could have formed in the early Universe independently of baryonic physics. In the standard Schwarzschild case, PBHs with initial masses $M>10^{17}g$ are typically considered viable dark matter candidates, as they would not have emitted enough Hawking radiation to be excluded by existing observational constraints. This translates into an upper bound on the fraction $f_{\text{PBH}}$ of dark matter that PBHs can constitute, depending on their mass. In our analysis, we have shown that regular black holes radiate less efficiently (figure 11) due to their lower temperature (figure 10), which qualitatively suggests that the lower bound on their allowed mass could be relaxed. As a consequence, the constraints on $M$ and $f_{\text{PBH}}$ are less stringent, potentially allowing for a larger fraction of dark matter to be composed of PBHs. This behavior is consistent with the recent findings [124, 125, 126], where it was demonstrated that other regular PBHs can evade gamma-ray constraints more easily and thus expand the viable dark matter mass window.
@@ Acknowledgments
We would like to acknowledge A. Zhidenko and O. Stashko for useful discussions. The work of A. S. was supported by University of Catania and INFN Sezione di Catania. A.S. is grateful for the hospitality of the Institute of Physics at Silesian University.